OpenStudy (anonymous):

How many different types of Indeterminate forms? What is the steps to solve each one? Why we need logarithmic function?

OpenStudy (lgbasallote):

there are many indeterminate forms...usually there are 6 used in calculus \[\infty - \infty\] \[1^\infty\] \[\frac 00\] \[\frac \infty \infty\] \[0 \times \infty\] \[\infty^0\] ^those are the most commonly used in calc

OpenStudy (anonymous):

thank you